# Robust Adaptation of Large Multimodal Models for Retrieval Augmented Hateful Meme Detection

### v2

@@ Abstract
Hateful memes have become a significant concern on the Internet, necessitating robust automated detection systems. While LMMs have shown promise in hateful meme detection, they face notable challenges like sub-optimal performance and limited out-of-domain generalization capabilities. Recent studies further reveal the limitations of both SFT and in-context learning when applied to LMMs in this setting. To address these issues, we propose a robust adaptation framework for hateful meme detection that enhances in-domain accuracy and cross-domain generalization while preserving the general vision-language capabilities of LMMs. Experiments on six meme classification datasets show that our approach achieves state-of-the-art performance, outperforming larger agentic systems. Moreover, our method generates higher-quality rationales for explaining hateful content compared to standard SFT, enhancing model interpretability.
@@ 1 Introduction
Figure 1: Comparison of rationales generated by SFT and RA-HMD Qwen2VL-7B models on the HatefulMemes dataset. The bar chart shows the winning rate of rationale quality based on pairwise comparisons between the two models. A more detailed analysis is provided in Appendix K.
@@
    RA-HMD demonstrates more robust out-of-domain generalization compared to SFT models. Notably, when RA-HMD is combined with a retrieval-augmented KNN classifier, it demonstrates state-of-the-art performance for out-of-domain meme classification. This setup leverages few-shot meme examples more effectively than in-context learning, addressing the challenge of adapting to rapidly evolving memes without retraining.
@@ 3.1 Preliminaries
[…]
@@ 3.2 RA-HMD Framework
Architecture enhancement Leveraging representations from large multimodal models (LMMs) for hateful meme classification is non-trivial, particularly when attempting to use LMM embeddings for classification while preserving the model’s original language generation capabilities. Prior work has explored various strategies for adapting these representations to classification and retrieval tasks. In our study, we similarly experimented with multiple adaptation methods. Appendix D provides a comprehensive summary of these efforts, including failure cases of previous approaches and key insights that ultimately guided the design of RA-HMD.
As illustrated in Figure 2, RA-HMD integrates an LMM with two additional trainable components: a Multilayer Perceptron (MLP) that projects the LMM final hidden state $\mathbf{h}_{i}$ into an embedding $\mathbf{g}_{i}$ for use in classification and retrieval; and an LRC operating on $\mathbf{g}_{i}$ Figure 2 shows how the architecture supports multiple fine-tuning and inference modes.
@@ 4.1 Comparing RA-HMD to Baseline Systems under Supervised Settings
Large Multimodal Models We experiment with three LMMs from two model families: LLaVA-1.5-7B Liu et al. (2023a), Qwen2VL-2B and Qwen2VL-7B Wang et al. (2024b). We report the performance of these LMMs in the following settings: pre-trained models with zero-shot and few-shot prompts using the LMH; SFT LMMs using the LMH; and classification using LRC under the RA-HMD fine-tuning framework. We further include the results with GPT-4o OpenAI (2024a) with optimized prompting for each dataset for reference. For GPT-4o, the token likelihood is not accessible to compute the AUC score.
Best Prior Models Visual Program Distillation (VPD) Hu et al. (2024) and ExplainHM Lin et al. (2024) are LLM agent-based systems. The remaining state-of-the-art models, including ISSUES Burbi et al. (2023), Pro-Cap Cao et al. (2023), RGCL Mei et al. (2024) and MemeCLIP Shah et al. (2024), are based on fine-tuning CLIP-based vision and language models. Detailed descriptions of these methods are provided in Appendix M.
@@
Across six datasets and three LMMs, fine-tuning with RA-HMD significantly improves performance over SFT (Table 1: #7, #8; #11, #12; #15, #16). Statistical significance tests comparing RA-HMD and SFT further validate these results, with all p-values below 0.05 (see Appendix C). Notably, as indicated in #16, Qwen2VL-7B fine-tuned with RA-HMD outperforms VPD-PaLI-X-55B on HatefulMemes. Moreover, RA-HMD improves upon RGCL with gains of over 4% in AUC and 3% in accuracy on HatefulMemes. These gains show RA-HMD’s effectiveness in improving LMMs for meme classification over SFT.
@@ 4.2 Comparing RA-HMD with Baseline Systems under Low-Resource Settings
We compare RA-HMD fine-tuned LMM with the RKC against the following systems: SFT LMMs with zero-shot and few-shot prompting using LMH; GPT-4o OpenAI (2024a); specialized low-resource systems (LOREHM Huang et al. (2024), Mod-hate Cao et al. (2024)). For GPT-4o, we report results without prompt optimization for each dataset, as this setting assumes the hate type is not known in advance. Further discussion and comparison of GPT-4o results can be found in Appendix J.
@@
Cross-domain fine-tuned LMMs show no consistent improvements over pre-trained LMMs for either zero-shot or few-shot prompting. Qwen2VL-7B zero-shot (#11 in Table 2) matches its SFT model performance (#13 in Table 1) on HatefulMemes and PrideMM but has performance degradation on the remaining four datasets.
@@
RA-HMD fine-tuned LMMs using RKC outperform the baseline SFT LMMs in both zero-shot and few-shot settings under the same cross-dataset fine-tuning settings. Notably, RA-HMD trained Qwen2VL-7B with RKC improves over the baseline SFT few-shot model by 21.6% in AUC and 19.3% in accuracy on HarMeme (Table 2 #11-13). The ablation study in Section 4.5, which varies the number of top k for RKC, further demonstrates that RKC uses demonstration examples more effectively than the few-shot in-context learning framework. Moreover, as shown in Appendix F, applying RKC to both pretrained and SFT LMMs results in worse performance compared to RA-HMD, underscoring the effectiveness of our fine-tuning strategy.
@@ 4.3 Effects of Two-Stage Fine-tuning
Table 3: Ablation study of RA-HMD two-stage fine-tuning framework on Qwen2VL-7B, evaluating the impact of Stage 1 and Stage 2 Fine-tuning. For $\mathcal{L}_{i}^{\textrm{Combined}}$, we jointly optimize the three loss objectives from both stages in a single training process as shown in Eq. 9.
@@ 4.4 Performance on General Vision-Language Benchmarks
Table 4 compares the pretrained Qwen2VL-2B with its SFT and RA-HMD variants, both fine-tuned on the HatefulMemes dataset, across three general vision-language benchmarks: MMMU Yue et al. (2023), SEED-Bench Li et al. (2023a), and GQA Hudson and Manning (2019). Evaluation settings and additional results are provided in Appendix I. The SFT model shows performance degradation across all three benchmarks, while RA-HMD maintains performance comparable to the pretrained model. These results indicate that RA-HMD robustly preserves the general vision-language capabilities of LMMs.
@@
Table 4: Comparison of the pretrained, SFT, and RA-HMD Qwen2VL-2B models on three general vision-language benchmarks. The SFT and RA-HMD models are fine-tuned on the HatefulMemes dataset.
@@ 4.8 Demonstration examples
In Appendix L, we present a case analysis comparing the classification results of RA-HMD and SFT. Appendix K provides example rationales generated by each model.
@@ 5 Conclusion
Training large-scale models is computationally intensive and contribute to global warming due to heavy GPU/TPU usage. However, our approach mitigates this issue by fine-tuning LMMs using quantized LoRA, a parameter-efficient method. As a result, our system can be trained in under four hours on a single consumer-grade GPU RTX3090, costing less than 1 USD, significantly reducing both training time and computational cost compared to full-scale LMM fine-tuning. Furthermore, since our method generalizes across different domains without requiring retraining, it further minimizes computational overhead.
@@ Appendix B Experiment Setup and Implementation Details
B.1 LLaVA and Qwen2VL experiments
@@
For Qwen2VL fine-tuning, we employ the officially recommended fine-tuning library LLaMA-Factory 0.9.1444https://github.com/hiyouga/LLaMA-Factory with official hyperparameter settings for downstream tasks in both the SFT and RA-HMD stage 1 fine-tuning. For few-shot learning with Qwen2VL, we follow the official multi-round conversation prompt format to ensure consistency with the model’s intended usage.
@@ Appendix D Insights for using LMMs representation for Meme Classification
In our experiments, we found that mean pooling generally underperformed compared to using the final token embedding. While this approach achieved a reasonable classification accuracy of about 77% on the HatefulMemes dataset using Qwen2VL-7B (RA-HMS Qwen2VL-7B has an accuracy of 82%), it had a major drawback: the language generation ability of the model was completely compromised. In practice, the model became unable to generate coherent text, suggesting that the learned representations were no longer aligned with the original language modeling objective.
@@ Appendix F Comparing Out-of-Domain Generalization Across Model Variants with In-Context Learning and RKC
  • RA-HMD-trained Qwen2VL-7B exhibits more robust generalization in the cross-dataset setting. Its in-context learning performance in few-shot scenarios consistently surpasses that of other models.
@@
Table 8: Comparing Pre-trained, SFT and RA-HMD systems with few-shot learning and RKC with Qwen2VL-7B under cross-dataset settings. See Table 2
@@ Appendix G Comparing Different Inference Modes
Table 9 compares Qwen2VL-7B fine-tuned with RA-HMD using the three classifiers. Our results indicate that under supervised settings, the differences among the three inference modes are minimal. However, under cross-dataset settings, there is a significant disparity in generalization performance. Notably, RKC outperforms both LMH and LRC, underscoring its superior effectiveness in handling out-of-domain examples.
@@
Table 9: Comparing different inference modes using RA-HMD fine-tuned Qwen2VL-7B. RKC shows much better out-of-domain generalization compared to other inference modes.
@@ Appendix H Ablation study on the loss function
Table 10: Ablation study of RA-HMD two-stage fine-tuning framework on Qwen2VL-7B, evaluating the impact of removing any of the loss objectives.
@@ Appendix I Evaluation on General Vision-Language Benchmarks
Table 11: Comparison of the pretrained, SFT, and RA-HMD Qwen2VL-2B models on three general vision-language benchmarks. The SFT and RA-HMD models are fine-tuned on the HatefulMemes dataset.
Appendix J GPT-4o Results and Prompts
@@
GPT-4o (Specific Prompt) - $71.3$ $72.9$ $79.4$ $63.1$ $64.5$ $58.3$ $58.1$ $75.3$ $73.7$
GPT-4o (General Prompt) - $66.40$ $68.4$ $72.90$ $55.39$ $55.11$ $61.07$ $51.07$ $63.79$ $62.28$
@@
Appendix K Rationale Generated by LMMs
We compare meme explanations generated by the SFT and RA-HMD fine-tuned Qwen2VL-7B models (Table 1 rows 15 and 16). RA-HMD is evaluated after Stage-1 fine-tuning, as Stage-2 does not fine-tune the Language Model Head for language generation. For evaluation, we prompt the above two models with:
@@
For reference, we include sample explanations of the meme examples in the Case Analysis in Appendix L Table 13.
@@
Table 13: Visualization of cases from SFT Qwen2VL-7B and RA-HMD Qwen2VL-7B Models on the HatefulMemes Dataset. Case 5 contains an insect in the meme; we applied a blurring filter to obscure it.
@@
Table 14: The error cases of SFT Qwen2VL-7B and RA-HMD Qwen2VL-7B models on HatefulMemes dataset
Appendix L Case Analysis
L.1 Comparing SFT and RA-HMD Predictions
Table 13 presents examples where our RA-HMD method successfully corrects prediction errors made by the SFT model on Qwen2VL-7B. Cases 1-4 involve hateful memes, while Cases 5-6 are benign memes that the SFT model misclassified, primarily due to poor multimodal alignment. These examples require a deep, joint understanding of both the image and text, a challenge that our RA-HMD effectively addresses. For example, in Case 2, the model needs to use its understanding of Japanese culture and associate this knowledge with the visual cues in the image.
L.2 Error Analysis
@@
Appendix M Baseline Methods

### v3

@@ Abstract
Hateful memes have become a significant concern on the Internet, necessitating robust automated detection systems. While Large Multimodal Models (LMMs) have shown promise in hateful meme detection, they face notable challenges like sub-optimal performance and limited out-of-domain generalization capabilities. Recent studies further reveal the limitations of both supervised fine-tuning (SFT) and in-context learning when applied to LMMs in this setting. To address these issues, we propose a robust adaptation framework for hateful meme detection that enhances in-domain accuracy and cross-domain generalization while preserving the general vision-language capabilities of LMMs. Analysis reveals that our approach achieves improved robustness under adversarial attacks compared to SFT models. Experiments on six meme classification datasets show that our approach achieves state-of-the-art performance, outperforming larger agentic systems. Moreover, our method generates higher-quality rationales for explaining hateful content compared to standard SFT, enhancing model interpretability. Code available at https://github.com/JingbiaoMei/RGCL
@@ 1 Introduction
Figure 1: Comparison of rationales generated by SFT and RA-HMD Qwen2-VL-7B models on the HatefulMemes dataset. The bar chart shows the winning rate of rationale quality based on pairwise comparisons between the two models. A more detailed analysis is provided in Appendix L.
@@
    RA-HMD demonstrates more robust out-of-domain generalization compared to SFT models. Notably, when RA-HMD is combined with a retrieval-augmented KNN classifier, it demonstrates state-of-the-art performance for out-of-domain meme classification. Moreover, this setup enhances robustness against adversarial attacks and leverages few-shot meme examples more effectively than in-context learning, thereby addressing the challenge of adapting to rapidly evolving memes without the need for retraining.
@@ 3.1 Preliminaries
Large Multimodal Models Some prior work has approached hateful meme detection via text generation with LMMs, where the LMM takes a meme $(I_{i},T_{i})$ as an input to predict a single token label $\hat{y}_{i}^{LMH}\in\{\textrm{``benign'', ``hateful''}\}$ Lin et al. (2024). We refer to the final linear layer of the LMM as the LM Head (LMH), which maps hidden representations to a probability distribution over the vocabulary via a softmax function. For meme classification, the LMH decodes the hidden state of the last token and generates the output label. This contrasts with approaches based on CLIP, which train Logistic Regression Classifiers (LRC) on encoder CLS tokens Kumar and Nandakumar (2022).
@@ 3.2 RA-HMD Framework
Architecture enhancement Leveraging representations from large multimodal models (LMMs) for hateful meme classification is non-trivial, particularly when attempting to use LMM embeddings for classification while preserving the model’s original language generation capabilities.
Prior work has explored various strategies for adapting these representations to classification and retrieval tasks. In our study, we similarly experimented with multiple adaptation methods. Appendix D provides a comprehensive summary of these efforts, including failure cases of previous approaches and key insights that ultimately guided the design of RA-HMD.
A central takeaway is that earlier adaptation methods enabled LMMs to perform retrieval but failed to preserve their ability to generate text simultaneously. In contrast, our proposed architecture, combined with a two-stage training procedure, successfully addresses this limitation.
As illustrated in Figure 2, RA-HMD integrates an LMM with two additional trainable components: a Multilayer Perceptron (MLP) that projects the LMM final hidden state $\mathbf{h}_{i}$ into an embedding $\mathbf{g}_{i}$ for use in classification and retrieval; and an LRC operating on $\mathbf{g}_{i}$ . Figure 2 shows how the architecture supports multiple fine-tuning and inference modes.
@@ 4.1 Comparing RA-HMD to Baseline Systems under Supervised Settings
Large Multimodal Models We experiment with three LMMs from two model families: LLaVA-1.5-7B Liu et al. (2023a), Qwen2-VL-2B and Qwen2-VL-7B Wang et al. (2024b). We report the performance of these LMMs in the following settings: pre-trained models with zero-shot and few-shot prompts using the LMH; SFT LMMs using the LMH; and classification using LRC under the RA-HMD fine-tuning framework. We further include the results with GPT-4o OpenAI (2024a) with optimized prompting for each dataset for reference. For GPT-4o, the token likelihood is not accessible to compute the AUC score.
Best Prior Models Visual Program Distillation (VPD) Hu et al. (2024) and ExplainHM Lin et al. (2024) are LLM agent-based systems. The remaining state-of-the-art models, including ISSUES Burbi et al. (2023), Pro-Cap Cao et al. (2023), RGCL Mei et al. (2024) and MemeCLIP Shah et al. (2024), are based on fine-tuning CLIP-based vision and language models. Detailed descriptions of these methods are provided in Appendix N.
@@
Across six datasets and three LMMs, fine-tuning with RA-HMD significantly improves performance over SFT (Table 1: #7, #8; #11, #12; #15, #16). Statistical significance tests comparing RA-HMD and SFT further validate these results, with all p-values below 0.05 (see Appendix C). Notably, as indicated in #16, Qwen2-VL-7B fine-tuned with RA-HMD outperforms VPD-PaLI-X-55B on HatefulMemes. Moreover, RA-HMD improves upon RGCL with gains of over 4% in AUC and 3% in accuracy on HatefulMemes. These gains show RA-HMD’s effectiveness in improving LMMs for meme classification over SFT.
@@ 4.2 Comparing RA-HMD with Baseline Systems under Low-Resource Settings
We compare RA-HMD fine-tuned LMM with the RKC against the following systems: SFT LMMs with zero-shot and few-shot prompting using LMH; GPT-4o OpenAI (2024a); specialized low-resource systems (LOREHM Huang et al. (2024), Mod-hate Cao et al. (2024)). For GPT-4o, we report results without prompt optimization for each dataset, as this setting assumes the hate type is not known in advance. Further discussion and comparison of GPT-4o results can be found in Appendix K.
@@
Cross-domain fine-tuned LMMs show no consistent improvements over pre-trained LMMs for either zero-shot or few-shot prompting. Qwen2-VL-7B zero-shot (#11 in Table 2) matches its SFT model performance (#13 in Table 1) on HatefulMemes and PrideMM but has performance degradation on the remaining four datasets.
@@
RA-HMD fine-tuned LMMs using RKC outperform the baseline SFT LMMs in both zero-shot and few-shot settings under the same cross-dataset fine-tuning settings. Notably, RA-HMD trained Qwen2-VL-7B with RKC improves over the baseline SFT few-shot model by 21.6% in AUC and 19.3% in accuracy on HarMeme (Table 2 #11-13). The ablation study in Section 4.6, which varies the number of top k for RKC, further demonstrates that RKC uses demonstration examples more effectively than the few-shot in-context learning framework. Moreover, as shown in Appendix F, applying RKC to both pretrained and SFT LMMs results in worse performance compared to RA-HMD, underscoring the effectiveness of our fine-tuning strategy.
@@ 4.3 Effects of Two-Stage Fine-tuning
Table 3: Ablation study of RA-HMD two-stage fine-tuning framework on Qwen2-VL-7B, evaluating the impact of Stage 1 and Stage 2 Fine-tuning. For $\mathcal{L}_{i}^{\textrm{Combined}}$, we jointly optimize the three loss objectives from both stages in a single training process as shown in Eq. 9.
@@ 4.4 Performance on General Vision-Language Benchmarks
Table 4 compares the pretrained Qwen2-VL-2B with its SFT and RA-HMD variants, both fine-tuned on the HatefulMemes dataset, across three general vision-language benchmarks: MMMU Yue et al. (2023), SEED-Bench Li et al. (2023a), and GQA Hudson and Manning (2019). Evaluation settings and additional results are provided in Appendix I. The SFT model shows performance degradation across all three benchmarks, while RA-HMD maintains performance comparable to the pretrained model. These results indicate that RA-HMD robustly preserves the general vision-language capabilities of LMMs.
@@
Table 4: Comparison of the pretrained, SFT, and RA-HMD Qwen2-VL-2B models on three general vision-language benchmarks. The SFT and RA-HMD models are fine-tuned on the HatefulMemes dataset.
4.5 Robustness Under Adversarial Attack
We follow Aggarwal et al. 2023 to assess the system’s robustness under adversarial attack. Specifically, we adopt the SaltPepper-I-High attack described in their work, which injects white and black pixels across the image in a manner that does not compromise the overall perception of semantic content.
We compare the SFT and the RA-HMD tuned Qwen2-VL-7B systems (corresponding to #15 and #16 in Table 1). As shown in Table 5, RA-HMD consistently outperforms SFT on SaltPepper-I-High–perturbed data, exhibiting less severe performance degradation. Moreover, when the perturbed examples are incorporated into the retrieval database alongside the original ones, robustness improves further. This shows that our RA-HMD retrieval-guided approach offers a simple yet effective way to enhance system robustness against adversarial attacks as soon as these attacks are detected.
HatefulMemes
Model AUC Acc.
Baseline (From Table 1)
SFT 86.3 78.6
RA-HMD + LRC 91.1 82.1
RA-HMD + RKC 90.8 81.8
Under Adversarial Attack
SFT 80.5 (-5.8) 72.3 (-6.3)
RA-HMD + LRC 84.4 (-6.7) 75.5 (-6.6)
RA-HMD + RKC 86.8 (-4.0) 76.6 (-5.2)
w/ Augmented DB 88.4 (-2.4) 78.4 (-3.4)
Table 5: Comparison of the SFT and RA-HMD Qwen2-VL-7B models on HatefulMemes under adversarial attack. Values in parentheses denote performance drop compared to non-attack.
@@ 4.8 Comparing Different Inference Modes
4.9 Comparing Rationales Generated by Models
We compare meme explanations generated by the SFT and RA-HMD fine-tuned Qwen2-VL-7B models (Table 1 rows 15 and 16). RA-HMD is evaluated after Stage-1 fine-tuning, as Stage-2 does not fine-tune the Language Model Head for language generation.
Following prior work Yang et al. (2023), we assess explanation quality using an LLM judge using pair-wise comparison and also a rubric-based evaluation. We evaluate the SFT and RA-HMD fine-tuned Qwen2-VL-7B models on the validation set of HatefulMemes dataset where the ground truth human annotated reasoning is available as a reference Hee et al. (2023). The LLM judge measures how closely model-generated explanations align with human rationales. Full details of the evaluation procedure are provided in Appendix L.
The pair-wise comparison results are:
  • RA-HMD beats SFT: 61.5
  • RA-HMD ties SFT: 13.8
  • SFT beats RA-HMD: 24.7
These results are visualized in Figure 1.
For the rubric-based evaluation (scored on a scale of 0–10), the SFT baseline achieved an average score of 6.6, while RA-HMD obtained a higher score of 7.5, further indicating stronger alignment with human rationales.
Based on the analysis of the generated explanations, we find that improvements in classification accuracy are supported by a deeper semantic understanding of memes. For challenging examples, where comprehension of background events or fine-grained details from the image is required, the RA–HMD–tuned system generates explanations that are both more accurate and more informative.
@@ 4.10 Demonstration examples
In Appendix M, we present a case analysis comparing the classification results of RA-HMD and SFT. Appendix L provides example rationales generated by each model.
@@ 5 Conclusion
Training large-scale models is computationally intensive and contributes to global warming due to heavy GPU usage. However, our approach mitigates this issue by fine-tuning LMMs using quantized LoRA, a parameter-efficient method. As a result, our system can be trained in under four hours on a single consumer-grade GPU RTX3090, costing less than 1 USD, significantly reducing both training time and computational cost compared to full-scale LMM fine-tuning. Furthermore, since our method generalizes across different domains without requiring retraining, it further minimizes computational overhead.
Acknowledgments
Jingbiao Mei is supported by Cambridge Commonwealth, European and International Trust for the undertaking of the PhD in Engineering at the University of Cambridge.
Jinghong Chen is supported by the Warwick Postgraduate Studentship from Christ’s College and the Huawei Hisilicon Studentship for the undertaking of the PhD in Engineering at the University of Cambridge.
Guangyu Yang is supported by Cambridge Commonwealth, European and International Trust for the undertaking of the PhD in Engineering at the University of Cambridge.
Weizhe Lin is supported by a Research Studentship funded by Toyota Motor Europe (RG92562(24020)) for the undertaking of the PhD in Engineering at the University of Cambridge.
Prof. Bill Byrne holds concurrent appointments as a Professor of Information Engineering at Cambridge University and as an Amazon Scholar. This publication describes work performed at Cambridge University and is not associated with Amazon.
We would also like to thank all the reviewers for their knowledgeable reviews that helped us strengthen the contribution.
@@ Appendix B Experiment Setup and Implementation Details
B.1 LLaVA and Qwen2-VL experiments
@@
For Qwen2-VL fine-tuning, we employ the officially recommended fine-tuning library LLaMA-Factory 0.9.1444https://github.com/hiyouga/LLaMA-Factory with official hyperparameter settings for downstream tasks in both the SFT and RA-HMD stage 1 fine-tuning. For few-shot learning with Qwen2-VL, we follow the official multi-round conversation prompt format to ensure consistency with the model’s intended usage.
@@ Appendix D Insights for using LMMs representation for Meme Classification
In our experiments, we found that mean pooling generally underperformed compared to using the final token embedding. While this approach achieved a reasonable classification accuracy of about 77% on the HatefulMemes dataset using Qwen2-VL-7B (RA-HMS Qwen2-VL-7B has an accuracy of 82%), it had a major drawback: the language generation ability of the model was completely compromised. In practice, the model became unable to generate coherent text, suggesting that the learned representations were no longer aligned with the original language modeling objective.
@@ Appendix F Comparing Out-of-Domain Generalization Across Model Variants with In-Context Learning and RKC
  • RA-HMD-trained Qwen2-VL-7B exhibits more robust generalization in the cross-dataset setting. Its in-context learning performance in few-shot scenarios consistently surpasses that of other models.
@@
Table 9: Comparing Pre-trained, SFT and RA-HMD systems with few-shot learning and RKC with Qwen2-VL-7B under cross-dataset settings. See Table 2
@@ Appendix G Comparing Different Inference Modes
Table 10 compares Qwen2-VL-7B fine-tuned with RA-HMD using the three classifiers. Our results indicate that under supervised settings, the differences among the three inference modes are minimal. However, under cross-dataset settings, there is a significant disparity in generalization performance. Notably, RKC outperforms both LMH and LRC, underscoring its superior effectiveness in handling out-of-domain examples.
@@
Table 10: Comparing different inference modes using RA-HMD fine-tuned Qwen2-VL-7B. RKC shows much better out-of-domain generalization compared to other inference modes.
@@ Appendix H Ablation study on the loss function
Table 11: Ablation study of RA-HMD two-stage fine-tuning framework on Qwen2-VL-7B, evaluating the impact of removing any of the loss objectives.
@@ Appendix I Evaluation on General Vision-Language Benchmarks
Table 12: Comparison of the pretrained, SFT, and RA-HMD Qwen2-VL-2B models on three general vision-language benchmarks. The SFT and RA-HMD models are fine-tuned on the HatefulMemes dataset.
Appendix J Task-specific fine-tuning and Multi-task fine-tuning
Here, we compare RA-HMD–tuned Qwen2-VL-7B models under two settings: (i) task-specific fine-tuning, where separate models are trained for each dataset, and (ii) multi-task fine-tuning, where a single model is jointly trained on all datasets. As shown in Table 13, multi-task fine-tuning results on Qwen2-VL-7B with RA-HMD suggest that a single system fine-tuned on all available datasets performs comparably to domain-specific systems.
Model HatefulMemes HarMeme MAMI Harm-P MultiOFF PrideMM
Single-task fine-tuning 82.1 88.1 79.9 91.6 71.1 78.1
Multi-task fine-tuning 82.0 87.8 80.4 91.6 70.4 78.5
Table 13: Accuracy comparison between single-task and multi-task fine-tuning across six datasets.
Appendix K GPT-4o Results and Prompts
@@
GPT-4o (Specific Prompt) - $71.3$ - $72.9$ - $79.4$ $63.1$ $64.5$ $58.3$ $58.1$ $75.3$ $73.7$
GPT-4o (General Prompt) - $66.40$ - $68.4$ - $72.90$ $55.39$ $55.11$ $61.07$ $51.07$ $63.79$ $62.28$
@@
Appendix L Rationale Generated by LMMs
We compare meme explanations generated by the SFT and RA-HMD fine-tuned Qwen2-VL-7B models (Table 1 rows 15 and 16). RA-HMD is evaluated after Stage-1 fine-tuning, as Stage-2 does not fine-tune the Language Model Head for language generation. For evaluation, we prompt the above two models with:
@@
Additionally, we conduct a rubric-based analysis, following the procedure in Yang et al. 2023. Specifically, we prompted the LLM judge to assign a score from 1 to 10 to each system’s rationale, based on how well it aligns with the human-written explanations. We assess the SFT and RA-HMD tuned Qwen2-VL-7B on the dev-seen split of the Facebook HatefulMemes dataset. Our results show that the SFT baseline received an average score of 6.6, while RA-HMD scored higher at 7.5, further indicating stronger alignment with human rationales.
For reference, we include sample explanations of the meme examples in the Case Analysis in Appendix M Table 15.
@@
Table 15: Visualization of cases from SFT Qwen2-VL-7B and RA-HMD Qwen2-VL-7B Models on the HatefulMemes Dataset. Case 5 contains an insect in the meme; we applied a blurring filter to obscure it. Furthermore, faces are also blurred.
@@
Table 16: The error cases of SFT Qwen2-VL-7B and RA-HMD Qwen2-VL-7B models on HatefulMemes dataset
Appendix M Case Analysis
M.1 Comparing SFT and RA-HMD Predictions
Table 15 presents examples where our RA-HMD method successfully corrects prediction errors made by the SFT model on Qwen2-VL-7B. Cases 1-4 involve hateful memes, while Cases 5-6 are benign memes that the SFT model misclassified, primarily due to poor multimodal alignment. These examples require a deep, joint understanding of both the image and text, a challenge that our RA-HMD effectively addresses. For example, in Case 2, the model needs to use its understanding of Japanese culture and associate this knowledge with the visual cues in the image.
M.2 Error Analysis
@@
Appendix N Baseline Methods